OpenStudy (anonymous):

The center of a circle is at (5, - 12) and it has a radius of 16. What is the equation of the circle?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: the general equation of any circle is (x-h)^2+(y-k)^2=r^2

jimthompson5910 (jim_thompson5910):

the center of that circle above is (h,k) the radius is r

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

(h,k) = (5,-12) so h = 5, k = -12

jimthompson5910 (jim_thompson5910):

the radius is 16, so r = 16

jimthompson5910 (jim_thompson5910):

plug h = 5, k = -12, r = 16 into (x-h)^2+(y-k)^2=r^2

OpenStudy (anonymous):

Do you mind working it out more with me please? @jim_thompson5910

jimthompson5910 (jim_thompson5910):

what do you get when you plug those values in?

OpenStudy (anonymous):

I got 33 :/

jimthompson5910 (jim_thompson5910):

you shouldn't replace x or y

OpenStudy (anonymous):

I mean 66

jimthompson5910 (jim_thompson5910):

(x-h)^2+(y-k)^2=r^2 (x-5)^2+(y-k)^2=r^2 ... plug in h = 5 (x-5)^2+(y-(-12))^2=r^2 ... plug in k = -12 (x-5)^2+(y-(-12))^2=4^2 ... plug in r = 4 and simplify from here

jimthompson5910 (jim_thompson5910):

you won't get a single numeric value because you aren't replacing x or y with values

jimthompson5910 (jim_thompson5910):

oh sry, typo, I meant to say on the last line (x-5)^2+(y-(-12))^2=16^2 since the radius is 16 and not 4

jimthompson5910 (jim_thompson5910):

anyways, you have to simplify (x-5)^2+(y-(-12))^2=16^2 to get the answer

OpenStudy (anonymous):

help me simplify this because im having a hard time and its not you your correct with everything I just need to see you simplify and fill in the x and y's in the equation

jimthompson5910 (jim_thompson5910):

what does y - (-12) simplify to

OpenStudy (anonymous):

what # replaces y?

jimthompson5910 (jim_thompson5910):

y is left alone

jimthompson5910 (jim_thompson5910):

y is just a variable and don't worry about it

jimthompson5910 (jim_thompson5910):

y - (-12) turns into ???

OpenStudy (anonymous):

I got 12

jimthompson5910 (jim_thompson5910):

where did y go?

OpenStudy (anonymous):

I put 0 for y

jimthompson5910 (jim_thompson5910):

just leave it as it is

jimthompson5910 (jim_thompson5910):

y - (-12) simplifies to y + 12

jimthompson5910 (jim_thompson5910):

because subtracting a negative is the same as adding a positive

jimthompson5910 (jim_thompson5910):

what is 16^2 ?

OpenStudy (anonymous):

256?

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (anonymous):

so my answer would be C?

jimthompson5910 (jim_thompson5910):

correct again

OpenStudy (anonymous):

Thank You So Much! do you think you can help me with this one?

jimthompson5910 (jim_thompson5910):

ok I can help with one more how far did you get with this one?

OpenStudy (anonymous):

I just started it

jimthompson5910 (jim_thompson5910):

any ideas on how to answer?

jimthompson5910 (jim_thompson5910):

If not, then I recommend finding the center first

OpenStudy (anonymous):

ok how do I do that

jimthompson5910 (jim_thompson5910):

the center is the midpoint of the segment from (2,0) to (2,-14)

OpenStudy (anonymous):

do I count the squares ?

jimthompson5910 (jim_thompson5910):

you could, or you can use the midpoint formula

OpenStudy (anonymous):

I counted 7 down the (2,0) to (2,-14)

jimthompson5910 (jim_thompson5910):

to what point?

OpenStudy (anonymous):

2 to 14

jimthompson5910 (jim_thompson5910):

you are correct in saying you go 7 units down

jimthompson5910 (jim_thompson5910):

since 7 is half of 14

jimthompson5910 (jim_thompson5910):

start at (2,0) travel 7 units down where do you end up?

OpenStudy (anonymous):

14

jimthompson5910 (jim_thompson5910):

that doesn't make any sense

jimthompson5910 (jim_thompson5910):

you have to end up at an ordered pair or point

OpenStudy (anonymous):

(2,-14)?

jimthompson5910 (jim_thompson5910):

that's the other endpoint

jimthompson5910 (jim_thompson5910):

but I want to know the center of this ellipse

jimthompson5910 (jim_thompson5910):